# High-Dimensional Independence Testing via Maximum and Average Distance Correlations

### v2

@@ 3.3 Limiting Null Distribution and Chi-square-based Tests
Assume that $X$ and $Y$ are independent, and that each dimension within $X$ and $Y$ is also independent. For sufficiently large $n$ and sufficiently small $\alpha$, it holds that
$n\cdot{\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\preceq_{\alpha}U-1,$
@@
Under the same condition in Theorem 4, and further assuming that both $n$ and $pq$ increase to infinity, it holds that
$n\sqrt{pq}\cdot{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,2).$
Note that this coincides with the limiting null distribution of the original distance correlation in high-dimensions (Szekely and Rizzo,, 2013). Moreover, while Theorem 4 holds for any $pq$, Theorem 5 requires $pq$ to increase. In practice, we have found that $pq>30$ suffices for a good approximation. Alternatively, if we assume each marginal distance correlation actually follows $\chi_{1}^{2}$, then the null distribution of $npq\cdot{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ equals $\chi_{pq}^{2}-pq$. This provides a better empirical approximation for small $pq$, and as $pq$ increases, it also converges to $N(0,2)$ after dividing by $\sqrt{pq}$.
@@
  • For the maximum distance correlation, we let $z=n\cdot{\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1$, and compute the p-value as
@@
  • For the average distance correlation, we let $z=n\cdot{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})+1$, and compute the p-value as
@@
Regarding the assumption of independence among dimensions within $\mathbf{X}$ and $\mathbf{Y}$, we shall clarify that even when this assumption is not met, the chi-square-based test remains approximately valid and is often more conservative. In other words, the p-value obtained through the chi-square-based test tends to be larger than the p-value obtained through the permutation test. For example, in the extreme scenario where $X^{1}=X^{2}=\ldots=X^{p}$ and $Y^{1}=Y^{2}=\ldots=Y^{q}$, Theorem 4 still holds, but with a conservative bound. In this case, a tighter bound for the null distribution should be $F_{\chi_{1}^{2}}^{1}(z)$.
Empirically, the presence of inter-dimension dependence appears to have relatively little impact on both tests as long as $n$ is moderate. However, for small samples, the p-value for the maximum distance correlation can be more conservative. In such cases, a permutation test may provide a more accurate result while remaining cost-effective for small $n$. These behaviors will be observed in the experiments section
@@ 4 Simulation Study
In our simulation study, we first demonstrate that the chi-square-based distribution provides an accurate approximation of the true null distribution. Next, we evaluate the testing power of the maximum and average tests on a variety of multivariate dependence.
@@ 4.1 Chi-Square vs True Null
Figure 1: Compare the chi-square distribution and true null distribution for distance correlation. The top row considers the Euclidean distance (DCor), while the bottom row employs the Gaussian kernel (HSIC). In the first column, we compare the null approximation for the original statistic as presented in Theorem 1. In the second and third columns, we compare the null approximation for the maximum and average statistics, respectively, based on Theorems 4 and 5.
@@ 4.2 Fixed $\delta(X,Y)$ with Increasing $pq$
We evaluate the testing power of original statistic, maximum statistic, and average statistic, in detecting multivariate dependence structures. The data is generated by sampling $X^{s}\sim Uniform(-1,1)$ for $s=1,\ldots,p$, using the $p\times 1$ vector $w=[1,\frac{1}{2},\frac{1}{3},\frac{1}{4},\frac{1}{5},0,\cdots,0]$, and considering
@@
  • Independence $(X,Y)$: $X^{s}\sim Uniform(-1,1)$ for $s=1,\ldots,p$ and $Y\sim Uniform(-1,1)$.
@@ 4.3 Increasing $\delta(X,Y)$ with Fixed $pq$
In this scenario, we examine different multivariate dependence structures where the number of marginally dependent dimension $\delta(X,Y)$ increases, while $pq$ remains fixed. Let $X^{s}\sim Uniform(-1,1)$ for $s=1,\ldots,p$, and consider
  • Linear $(X,Y)$: $Y^{s}=X^{s}$ for each $s\leq d$, and $Y^{s}\sim Uniform(-1,1)$ otherwise.
  • Trigonometry $(X,Y)$: $Y^{s}=sin(2\pi X^{s})$ for each $s\leq d$, and $Y^{s}\sim Uniform(-1,1)$ otherwise.
@@ 5 Real Data
This experiment aimed to investigate the presence of any dependency between the abundance levels of peptides in human plasma and the occurrence of cancers. Selected Reaction Monitoring (SRM) was employed as a targeted quantitative proteomics technique for measuring protein and peptide abundance in complex biological samples (Wang et al.,, 2011). A prior study utilized SRM to identify a total of $318$ peptides from a total of $98$ individuals, among whom $33$ were normal subjects, $10$ had pancreatic cancer, $24$ had colorectal cancer, and $28$ had ovarian cancer (Wang et al.,, 2017). Consequently, $\mathbf{X}$ represents the sample peptide levels with $p=318$. The data is publicly available on a GitHub repository in MATLAB format [^1] https://github.com/neurodata/MGC-paper/blob/master/Data/Preprocessed/proteomics.mat.
@@
We performed independence tests based on various combinations, including utilizing the entire sample dataset, where $\mathbf{Y}$ represents a label vector indicating the cancer type each subject has. Other test scenarios included: distinguishing normal individuals from others, where $\mathbf{Y}$ is a label vector with normal subjects as $1$ and all others as $2$; distinguishing pancreatic from colorectal cancer, with $\mathbf{Y}$ as a label vector where pancreatic subjects are labeled as $1$, colorectal subjects as $2$, and others as unused; and so forth.
@@
Our previous study (Vogelstein et al.,, 2019) indicated that all such testing combinations should yield significant p-value. Table 1 presents the test statistics and p-values for the maximum distance correlation, average distance correlation, and original distance correlation. While the original distance correlation performed well, there were three combinations where it failed to detect dependence: pancreatic vs others, colorectal vs others, and pancreatic vs colorectal. This may not be surprising, given that there were only $10$ subjects with pancreatic cancer, which is the smallest group in the dataset, and colorectal cancer is the second smallest group. Since the sample size $n$ is small in these cases, we also conducted permutation tests using $100$ random permutations for these insignificant pairs (results reported in brackets), and the results remained insignificant.
@@
The maximum correlation yielded significant results overall, except in three testing combinations involving pancreatic cancer. This aligns with the theoretical findings where the maximum correlation may not perform well for small sample sizes due to bias from the null distribution, and the chi-square test can be overly conservative. Therefore, we conducted permutation tests in these cases, and the results improved significantly, with all of them becoming statistically significant at type 1 error level $\alpha=0.07$.
@@
The average correlation yielded significant p-values in almost all combinations, as it is not sensitive to small sample sizes. The only exception was the test for pancreatic vs. colorectal cancer, where it did not yield a significant result, while the maximum method would have tested significant at $\alpha=0.07$. Its underperformance in this case suggests that $\delta(X,Y)$, the number of marginally dependent dimensions, could be very small, leading to its lack of sensitivity.
@@
Testing Combination n ${\sc\texttt{Dcor}}_{n}$ p-val ${\sc\texttt{MDcor}}_{n}$ p-val ${\sc\texttt{ADcor}}_{n}$ p-val
@@
PANC vs Others 98 $0.007$ 0.19(0.19)0.190.190.19(0.19)0.19 ( 0.19 ) $0.11$ $0.18(0.07)$ $0.003$ $0.01$
COLO vs Others 98 $0.022$ 0.07(0.07)0.070.070.07(0.07)0.07 ( 0.07 ) $0.15$ $0.02$ $0.013$ $<0.001$
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Science Foundation awards DMS-1921310 and DMS-2113099, the University of Delaware Data Science Institute Seed Funding Grant, and the Defense Advanced Research Projects Agency’s L2M program FA8650-18-2-7834.
@@ Acknowledgments
7 All Proofs
Theorem 2.
Given $\delta(X,Y)=0$, the average distance correlation satisfies:
${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
regardless of $pq$.
On the other hand, the maximum distance correlation satisfies:
${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0$
when $pq=o(e^{\sqrt{n}})$.
@@ 7 All Proofs
${\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0.$
@@
$Prob({\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})<\epsilon)$
$\geq Prob(n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})<n\epsilon)^{pq}$
@@
Here, the second line follows from basic order statistics, the third line follows from the existing dominance results by Theorem 1, the fourth line is based on the approximation of the standard normal distribution at the tail, where $x=(n\epsilon+1)^{0.5}=O(n^{0.5})$.
Therefore, it suffices to consider when $pq$ is a function of $n$, where we have:
$\lim_{n\rightarrow\infty}(1-\frac{e^{-n}}{\sqrt{n}})^{pq}\rightarrow 1,$
or equivalently:
$\lim_{n\rightarrow\infty}pq\log(1-\frac{1}{\sqrt{n}e^{n}})\rightarrow 0.$
By computing the limit while treating $pq$ as a function of $n$, and then using L’Hôpital’s rule (details omitted), we find that the above limit is 00 as long as:
$(pq)^{2}/(pq)^{\prime}=o(e^{n}/\sqrt{n}),$
where $(pq)^{\prime}$ represents its derivative with respect to $n$. A sufficient condition is therefore $pq=o(e^{\sqrt{n}})$. ∎
Theorem 3.
When $\delta(X,Y)>0$, the maximum distance correlation satisfies:
${\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})>0$
regardless of $pq$.
On the other hand, the average distance correlation satisfies:
${\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c\geq 0,$
@@
${\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c>0.$
Therefore, the maximum distance correlation is always greater than 00.
@@
$|{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})|$ $=|\sum\limits_{s\in[p],t\in[q]}{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})|/pq$
@@
Theorem 4.
Assume that $X$ and $Y$ are independent, and that each dimension within $X$ and $Y$ is also independent. For sufficiently large $n$ and sufficiently small $\alpha$, it holds that
$n\cdot{\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\preceq_{\alpha}U-1,$
where $U\sim F_{\chi_{1}^{2}}^{pq}(x)$.
@@
For each $s\in[p],t\in[q]$, we apply the upper-tail dominance property from Theorem 1 to each marginal distance correlation, i.e.,
$n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ $\preceq_{\alpha}\chi_{1}^{2}-1$ (1)
@@
$F_{n\cdot{\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1}(z)$
$=$ $\prod_{s\in[p],t\in[q]}Prob(n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})+1\leq z)$
@@
Consequently, we have $n\cdot{\sc\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1\preceq_{\alpha}U$. ∎
Theorem 5.
Under the same condition in Theorem 4, and further assuming that both $n$ and $pq$ increase to infinity, it holds that
$n\sqrt{pq}\cdot{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,2).$
@@
From Theorem 1, the limiting null distribution of each marginal distance correlation satisfies
$n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ $\stackrel{{\scriptstyle D}}{{\rightarrow}}\sum\limits_{i,j=1}^{\infty}w_{ij}(\mathcal{N}_{ij}^{2}-1),$
where the weights satisfy $w_{ij}\in[0,1]$ and $\sum\limits_{i,j=1}^{\infty}w_{ij}^{2}=1$, and $\mathcal{N}_{ij}$ are independent standard normal distribution. Therefore, as $n\rightarrow\infty$, each marginal distance correlation has an expected value of 00 and a variance of $2/n^{2}$.
By central limit theorem,
$n\sqrt{pq}\cdot{\sc\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$
$=$ $\sum\limits_{s\in[p],t\in[q]}(n\cdot{\sc\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t}))/\sqrt{pq}$
$\stackrel{{\scriptstyle D}}{{\rightarrow}}$ $N(0,2).$
Theorem 6.

### v3

@@ 3.3 Limiting Null Distribution
Assume that $X$ and $Y$ are independent, and that the coordinates of $X$ and $Y$ are independent within each variable. For sufficiently large $n$ and sufficiently small $\alpha$, it holds that
$n\cdot{\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})$ $\preceq_{\alpha}U-1,$
@@
Assume that $X$ and $Y$ are independent, and that the coordinates of $X$ and $Y$ are independent within each variable. As both $n$ and $pq$ increase to infinity, it holds that
$n\sqrt{pq}\cdot{\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,2).$
Note that this coincides with the limiting null distribution of the original distance correlation in high-dimensions [37]. Moreover, while Theorem 3 holds for any $pq$, Theorem 4 requires $pq$ to increase. In practice, we have found that $pq>30$ suffices for a good approximation. Alternatively, if we assume each marginal distance correlation actually follows $\chi_{1}^{2}$, then the null distribution of $npq\cdot{\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})$ equals $\chi_{pq}^{2}-pq$. This provides a better empirical approximation for small $pq$, and as $pq$ increases, it also converges to $N(0,2)$ after dividing by $\sqrt{pq}$.
Regarding the assumptions, we require that the coordinates of $X$ and $Y$ are independent. While this may appear restrictive at first glance, it can be further relaxed. Specifically, if we additionally assume that both $p$ and $q$ tend to infinity, the independence assumption can be relaxed to exchangeability, similar to Theorem 1 and Corollary 2 in [37]. This can be justified by de Finetti’s theorem, which states that an infinite exchangeable sequence is conditionally i.i.d. Our proofs remain valid under this setting, as the maximum or average distance correlation retains the same asymptotic distribution when conditioning on the latent variable.
3.4 Chi-square-based Tests
@@
  • For the maximum distance correlation, we let $z=n\cdot{\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1$, and compute the p-value as
@@
  • For the average distance correlation, we let $z=n\cdot{\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})+1$, and compute the p-value as
@@
3.5 Computational Complexity
Taken together, the maximum and average distance correlation statistics have a computational complexity of $O(pq\,n\log n)$. This is because each marginal distance correlation can be computed in $\mathcal{O}(n\log n)$ time under the Euclidean distance metric, thanks to [33, 34]. The chi-square-based tests incur no additional cost, so the overall complexity for testing remains $O(pq\,n\log n)$. However, if a permutation test with $R$ permutations is used, the total complexity increases to $O(R\,pq\,n\log n)$, and in practice, $R$ is typically 100 or more.
If one instead uses a non-Euclidean metric such as the Gaussian kernel, fast univariate algorithms are no longer available. In this case, the computation of the test statistic has a complexity of $O(pq\,n^{2})$, and using a permutation test raises the total complexity to $O(R\,pq\,n^{2})$.
For comparison, the original (non-marginal) distance correlation or kernel-based methods require $O((p+q)\,n^{2})$ time. However, since the pairwise distance or kernel matrix is computed only twice, the cost is typically treated as $O(n^{2})$ in practice. Therefore, marginal-based methods can be slower in terms of raw computation. Namely, in high-dimensional testing scenarios, potential gain in statistical power shall come at the cost of increased computational burden.
As an additional note, our average distance correlation performs similarly to the average distance covariance proposed in [25], since distance correlation is simply a normalized version of distance covariance, and their testing power has been shown to coincide [2]. However, our method offers a computational advantage: the average distance covariance typically requires a permutation test (costing $O(R\,pq\,n\log n)$), or alternatively, a transformation into a $t$-distribution for avoiding permutations. The latter approach requires computing marginal distance variances, resulting in a complexity of $O((pq+p^{2}+q^{2})\,n\log n)$, which can exceed the cost of a permutation test when $p>>q$ or vice versa.
In contrast, our average distance correlation, paired with the efficient chi-square test, achieves similar testing power and significantly lower computational cost, making it more suitable for large sample sizes and high-dimensional settings.
@@ 4 Simulation Study
In our simulation study, we first demonstrate that the chi-square-based distribution provides an accurate approximation of the true null distribution. Then we evaluate the testing power of the maximum and average tests on a variety of multivariate dependence.
@@ 4.1 Chi-Square vs True Null
Figure 1: Compare the chi-square distribution and true null distribution for distance correlation. The top row considers the Euclidean distance (DCor), while the bottom row employs the Gaussian kernel (HSIC). In the first column, we compare the null approximation for the original statistic. In the second and third columns, we compare the null approximation for the maximum and average statistics, respectively, based on Theorems 3 and 4.
@@ 4.2 Fixed $\delta(X,Y)$ with Increasing $pq$
We evaluate the testing power of original statistic, maximum statistic, and average statistic, in detecting multivariate dependence structures. The data is generated by sampling $X^{s}\sim\mbox{Uniform}(-1,1)$ for $s=1,\ldots,p$, using the $p\times 1$ vector $w=[1,\frac{1}{2},\frac{1}{3},\frac{1}{4},\frac{1}{5},0,\cdots,0]$, and considering
@@
  • Independence $(X,Y)$: $X^{s}\sim\mbox{Uniform}(-1,1)$ for $s=1,\ldots,p$ and $Y\sim\mbox{Uniform}(-1,1)$.
Here, each sample in $\mathbf{X}$ is independently and identically distributed (i.i.d.) from the distribution of $X$, and $\mathbf{Y}$ is constructed as a function of each corresponding sample in $\mathbf{X}$, except in the case of independence, where each sample in $\mathbf{Y}$ is also independently generated.
@@ 4.3 Increasing $\delta(X,Y)$ with Fixed $pq$
In this scenario, we examine different multivariate dependence structures where the number of marginally dependent dimension $\delta(X,Y)$ increases, while $pq$ remains fixed. Let $X^{s}\sim\mbox{Uniform}(-1,1)$ for $s=1,\ldots,p$, and consider
  • Linear $(X,Y)$: $Y^{s}=X^{s}$ for each $s\leq d$, and $Y^{s}\sim\mbox{Uniform}(-1,1)$ otherwise.
  • Trigonometry $(X,Y)$: $Y^{s}=sin(2\pi X^{s})$ for each $s\leq d$, and $Y^{s}\sim\mbox{Uniform}(-1,1)$ otherwise.
@@ 5 Real Data
This experiment aimed to investigate the presence of any dependency between the abundance levels of peptides in human plasma and the occurrence of cancers. Selected Reaction Monitoring (SRM) was employed as a targeted quantitative proteomics technique for measuring protein and peptide abundance in complex biological samples [38]. A prior study utilized SRM to identify a total of $318$ peptides from a total of $98$ individuals, among whom $33$ were normal subjects, $10$ had pancreatic cancer, $24$ had colorectal cancer, and $28$ had ovarian cancer [39]. Consequently, $\mathbf{X}$ represents the sample peptide levels with $p=318$. The data is publicly available on a GitHub repository in MATLAB format [^1] https://github.com/neurodata/MGC-paper/blob/master/Data/Preprocessed/proteomics.mat.
@@
We performed independence tests based on various combinations, including utilizing the entire sample dataset, where $\mathbf{Y}$ represents a label vector indicating the cancer type each subject has. Other test scenarios included: distinguishing normal individuals from others, where $\mathbf{Y}$ is a label vector with normal subjects as 00 and all others as $1$; distinguishing pancreatic from colorectal cancer, with $\mathbf{Y}$ as a label vector where pancreatic subjects are labeled as 00, colorectal subjects as $1$, and others as unused; and so forth. Note that categorical response variables can be directly incorporated in dependence testing. Binary variables, for instance, can be numerically encoded (e.g., as 0 and 1), while multi-category variables can be handled via one-hot encoding [40, 17].
@@
A previous study [41] indicated that all such testing combinations should yield significant p-value. Table 1 presents the test statistics and p-values for the maximum distance correlation, average distance correlation, and original distance correlation. While the original distance correlation performed well, there were three combinations where it failed to detect dependence: pancreatic vs others, colorectal vs others, and pancreatic vs colorectal. This may not be surprising, given that there were only $10$ subjects with pancreatic cancer, which is the smallest group in the dataset, and colorectal cancer is the second smallest group.
@@
The maximum correlation yielded significant results overall, except in three testing combinations involving pancreatic cancer. As the chi-square test for the maximum correlation may be overly conservation for small sample sizes, in these three cases, we further conducted permutation tests using $100$ random permutations (results reported in brackets), and the results improved significantly.
@@
The average correlation yielded significant p-values in almost all combinations, as its chi-square test is robust against small sample sizes. The only exception was the test for pancreatic vs. colorectal cancer, where it did not yield a significant result, while the maximum method is very close to significance. Its underperformance in this case suggests that $\delta(X,Y)$, the number of marginally dependent dimensions, could be very small, leading to its lack of sensitivity.
@@
Combination n ${\texttt{Dcor}}_{n}$ p-val ${\texttt{MDcor}}_{n}$ p-val ${\texttt{ADcor}}_{n}$ p-val
@@
PANC vs Others 98 $0.007$ $0.19$ $0.11$ $0.18(0.07)$ $0.003$ $0.01$
COLO vs Others 98 $0.022$ $0.07$ $0.15$ $0.02$ $0.013$ $<0.001$
@@ 6 Conclusion
Beyond their theoretical appeal, these methods are highly applicable to real-world problems, particularly in biomedical and biological sciences. In such fields, researchers are often interested in detecting complex, nonlinear, and high-dimensional dependencies, such as between gene expression profiles and disease subtypes, or between multi-omics data and clinical outcomes. Our methods are well suited to these tasks, as they operate in a fully nonparametric, model-free framework, require minimal assumptions, and are robust to data heterogeneity.
Moreover, the ability to decompose dependence through marginal statistics (and aggregate them meaningfully) offers practical interpretability and flexibility for applied scientists. Compared to classical approaches rooted in covariance-based inference (e.g., the Wishart distribution under multivariate normality), our approach targets a broader class of multivariate dependence structures.
Taken together, we believe the proposed distance correlation framework offers a valuable statistical tool for both theoretical development and applied data analysis.
Declarations
  • Availability of data and materials: The datasets generated and/or analysed during the current study are available in the Github repository, [https://github.com/neurodata/mgc-matlab].
  • Author Contributions: C.S.: conceptualization, method, theory, experiments, writing; Y.D.: conceptualization, theory, discussion. All authors have read and agreed to the published version of the manuscript.
  • Funding: This work was supported in part by the National Science Foundation DMS-1921310 and DMS-2113099, the University of Delaware Data Science Institute Seed Funding Grant, and the Defense Advanced Research Projects Agency’s L2M program FA8650-18-2-7834.
  • Acknowledgments: We sincerely thank the editor and anonymous reviewers for providing valuable and timely feedback that significantly improved the paper.
@@
A1 Technical Prerequisite
We first cite the main theorem of [35], which is necessary for subsequent proofs. It holds under the assumption of independence between $X$ and $Y$, finite-moments of $F_{XY}$, an increasing sample size $n$, and regardless of $p,q$ nor the specific distribution of $F_{XY}$.
Theorem A1.
For sufficiently large $n$, there exists $\alpha>0$ such that
$\mathcal{N}(0,2)\preceq_{\alpha}n\cdot{\texttt{Dcor}}_{n}(\mathbf{X},\mathbf{Y})\preceq_{\alpha}\chi_{1}^{2}-1$
regardless of the metric choice or marginal distributions.
Definition B2.
$F_{V}(z)\geq F_{U}(z)$
A2 Theorem Proofs
See 1
@@
${\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}0.$
@@
$Prob({\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})<\epsilon)$
$\geq Prob(n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})<n\epsilon)^{pq}$
@@
Here, the second line follows from basic order statistics, and the third line follows from the existing dominance results by Theorem A1.
Next, consider the asymptotic expansion of the standard normal tail probability:
$1-\Phi(x)=\frac{1}{\sqrt{2\pi}x}e^{-x^{2}/2}\left(1-\frac{1}{x^{2}}+\frac{3}{x^{4}}-\cdots\right),\quad\text{as }x\to\infty,$
where $\Phi(x)$ is the standard normal CDF. Applying this to our setting, where
$F_{\chi_{1}^{2}-1}^{pq}(n\epsilon)=\left[2\Phi(x)-1\right]^{pq},$
we immediately have
$F_{\chi_{1}^{2}-1}^{pq}(n\epsilon)=\left(1-\frac{2}{\sqrt{2\pi}x}e^{-x^{2}/2}+o\left(\frac{e^{-x^{2}/2}}{x^{3}}\right)\right)^{pq}=(1-\delta_{n})^{pq},$
where $x=\sqrt{n\epsilon+1}=O(\sqrt{n})$, which increases to infinity at any fixed $\epsilon$.
We aim to show $(1-\delta_{n})^{pq}\rightarrow 1$. It suffices for $pq\delta_{n}\rightarrow 0$. This is because:
$(1-\delta_{n})^{pq}=\exp\left(pq\cdot\log(1-\delta_{n})\right).$
As $\delta_{n}\to 0$, we use the Taylor expansion $\log(1-\delta_{n})=-\delta_{n}+o(\delta_{n})$, so
$(1-\delta_{n})^{pq}=\exp\left(-pq\cdot\delta_{n}+o(pq\cdot\delta_{n})\right).$
If $pq\cdot\delta_{n}\to 0$, then the exponent tends to zero, and hence
$(1-\delta_{n})^{pq}\to 1.$
To ensure $pq\delta_{n}\rightarrow 0$, a sufficient condition would be
$pq\cdot\frac{e^{-x^{2}/2}}{x}\to 0.$
To see why, if the above holds, then
$pq\cdot\frac{2}{\sqrt{2\pi}x}e^{-x^{2}/2}\to 0\quad\text{and}\quad pq\cdot o\left(\frac{e^{-x^{2}/2}}{x^{3}}\right)\to 0,$
and hence,
$pq\cdot\left(\frac{2}{\sqrt{2\pi}x}e^{-x^{2}/2}+o\left(\frac{e^{-x^{2}/2}}{x^{3}}\right)\right)\to 0.$
Finally, to satisfy
$pq\cdot\frac{e^{-x^{2}/2}}{x}\to 0$
with $x=\sqrt{n\epsilon+1}=O(\sqrt{n})$, we observe that
$\frac{e^{-x^{2}/2}}{x}=O\left(\frac{e^{-n/2}}{\sqrt{n}}\right).$
Therefore, a sufficient condition is
$pq=o\left(\sqrt{n}\cdot e^{n}\right).$
See 2
@@
${\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}c>0.$
Therefore, the maximum distance correlation shall converge to a positive constant.
@@
$|{\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})|$ $=|\sum\limits_{s\in[p],t\in[q]}{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})|/pq$
@@
See 3
@@
For each $s\in[p],t\in[q]$, we apply the upper-tail dominance property from Theorem A1 to each marginal distance correlation, i.e.,
$n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ $\preceq_{\alpha}\chi_{1}^{2}-1$
@@
$F_{n\cdot{\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1}(z)$
$=$ $\prod_{s\in[p],t\in[q]}Prob(n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})+1\leq z)$
@@
Consequently, we have $n\cdot{\texttt{MDcor}}_{n}(\mathbf{X},\mathbf{Y})+1\preceq_{\alpha}U$. ∎
See 4
@@
First,
$n\sqrt{pq}\cdot{\texttt{ADcor}}_{n}(\mathbf{X},\mathbf{Y})=\frac{1}{\sqrt{pq}}\sum_{s=1}^{p}\sum_{t=1}^{q}n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t}).$
As mentioned in the background section, the limiting null distribution of each marginal distance correlation satisfies
$n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ $\stackrel{{\scriptstyle D}}{{\rightarrow}}\sum\limits_{i,j=1}^{\infty}w_{ij}(\mathcal{N}_{ij}^{2}-1),$
where the weights satisfy $w_{ij}\in[0,1]$ and $\sum\limits_{i,j=1}^{\infty}w_{ij}^{2}=1$, and $\mathcal{N}_{ij}$ are independent standard normal distribution. Therefore, as $n\rightarrow\infty$, each $n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ has an expected value of 00 and a variance of
$\operatorname{Var}\left(\sum_{i,j}w_{ij}(\mathcal{N}_{ij}^{2}-1)\right)=2\sum_{i,j}(w_{ij})^{2}=2.$
Furthermore, since this limit is a (centered) quadratic form in independent sub-Gaussian random variables, it satisfies uniform exponential tail bounds. Specifically, there exists a constant $C>0$ such that
$\sup_{s,t,n}\mathbb{E}\left[\exp\left(C\cdot\left|n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\right|\right)\right]<\infty.$
Under the given assumptions, the marginal statistics $n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})$ are independent across $s\in[p]$, $t\in[q]$. Then we may view the full sum as a triangular array:
$Z_{n}:=\frac{1}{\sqrt{pq}}\sum_{s=1}^{p}\sum_{t=1}^{q}\left(n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\right).$
Each summand has mean zero and variance 2, so the total variance of $Z_{n}$ is
$\operatorname{Var}(Z_{n})=\frac{1}{pq}\cdot pq\cdot 2=2.$
To apply the triangular array Central Limit Theorem (e.g., Lindeberg-Feller CLT), it suffices to verify the Lindeberg condition:
$\forall\varepsilon>0,\quad\lim_{n\to\infty}\frac{1}{pq}\sum_{s,t}\mathbb{E}\left[\left(n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\right)^{2}\cdot\mathbf{1}\left\{\left|n\cdot{\texttt{Dcor}}_{n}(\mathbf{X}^{s},\mathbf{Y}^{t})\right|>\varepsilon\sqrt{pq}\right\}\right]=0.$
This condition is satisfied due to the exponential tail bounds established above, which ensure that large deviations decay faster than any polynomial in $pq$. Therefore, the contribution of rare large terms becomes negligible as $pq\to\infty$. Hence, the Lindeberg CLT applies, and we conclude:
[…]
See 5